OpenStudy (anonymous):

Solve, using linear combination: 4x + y = 5 3x + y = 3 A: (0, –3) B: (0, 5) C: (2, –3) D: (–3, 7)

OpenStudy (triciaal):

eqn 1 subtract eqn 2 to eliminate y solve for x substitute in either eqn to find y

OpenStudy (triciaal):

x = 2 option c check y anyway 4(x) + y =5 4(2) +y =5 8 + y =5 y =

OpenStudy (anonymous):

ok thank you :)

OpenStudy (triciaal):

welcome